equation of line which passes through the points (2,5) and (3,9)
Answer by god2012(113)

Solution:
Using 2 point form a line: (y2-y1)/(x2-x1) = (y-y1)/(x-x1)
(x1,y1) = (2, 5)
(x2,y2) = (3, 9)
(9-5)/(3-2) = (y-5)/(x-2) = 4
y-5 = 4x - 8
4x - y - 3 = 0